OpenStudy (anonymous):

I am trying to solve this problem 12√(-8)^12 I mutiplied 8 12 time then square that off with 12

OpenStudy (anonymous):

-8^12 = 16777216 \[\sqrt{16777216}\] = 4096 4096 x 12 = 49152

OpenStudy (anonymous):

Thanks, but how did you get 16777216

OpenStudy (m):

you get it by raising (-8) to the 12th power

OpenStudy (m):

anything you raise to even number exponent you get positive number

OpenStudy (anonymous):

\[\sqrt{5}(6\sqrt{2}-6\sqrt{25}\]

OpenStudy (anonymous):

I took what inside of the parenthesis and square and then substract